OpenStudy (anonymous):

1) £35-10% 2) £6-20% 3) £8.50-30% 4) £60-40% 5) £19-25% 6) £12-15%

OpenStudy (anonymous):

What do u want???

OpenStudy (anonymous):

the answers cause, in my maths class we learnt it but then i forgot,

OpenStudy (anonymous):

1) £35-10% = 35 - (10/100 * 35) = 35 - 3.5 = 31.5 2) £6-20% = 6 - (20/100 * 6) = 6 - 1.2 = 4.8

OpenStudy (anonymous):

u can solve the rest in the same way....☺

OpenStudy (anonymous):

thanks alot